# HydraLoRA: An Asymmetric LoRA Architecture for Efficient Fine-Tuning

### v1

@@ 3.2.3 Inference
During inference, HydraLoRA merges adapters by enabling routing computation based on the input. Specifically, since matrices B operate as linear functions, we initially compute a weighted average of the experts. Following this, we apply a PEFT transformation using the combined expertise.
The HydraLoRA significantly enhances training efficiency through an extremely parameter-efficient MoE formulation. Additionally, the intrinsic structural modularity of HydraLoRA facilitates rapid recovery and merging of the trained parameters during inference, leading to substantial memory savings.
4 Experiments
In this section, we detail our principal experiments. We begin with an overview of the experimental setup and implementation intricacies. Following this, we share our findings and offer a succinct interpretation.
4.1 Experiment Setting
Benchmarks
To explore the properties and commonalities of the LoRA asymmetric structure, we conduct experiments on both single and multiple domains to evaluate the effectiveness of HydraLoRA for profiling intrinsic components.
$\bullet$ Single domain.
  • General: we fine-tune with the general instruction tuning databricks-dolly-15k [DollyV2] for generic language capability and evaluate with MMLU [mmlu].
  • Medical: we fine-tune with GenMedGPT and clinic-10k from ChatDoctor [chatdoctor] for medicine applications and evaluate medical tasks in MMLU including three related tasks: “clinical knowledge”, “professional medicine” and “college medicine”.
  • Law: we fine-tune with two legal instruction tuning datasets Lawyer-Instruct [Lawyer-Instruct] and US-Terms [legal_lama] then evaluate with law tasks in MMLU including two related tasks: “professional law” and “international law”.
  • Math: we fine-tune with the training split of GSM8K [gsm8k] for mathematical reasoning and evaluate with test set of GSM8K.
  • Code: we fine-tune with CodeAlpaca [codealpaca] for code generation and evaluate with HumanEval [Humaneval].
$\bullet$ Multi-task domain.
  • Flanv2: We select a portion of the Flanv2 [flanv2] datasets covering Natural Language Understanding (NLU) and Natural Language Generation (NLG), which can be grouped into 10 distinct task clusters. Then we evaluate it with the Big-Bench Hard (BBH) benchmark. A detailed description of the benchmarks can be found in Appendix A.
Baselines
First, we compare HydraLoRA with different PEFT methods on single datasets:
  • Prompt Tuning [LesterAC21] adds task-specific prompts to the input, and these prompt parameters are updated independently of the pretrained model parameters which are frozen.
  • P-Tuning [P-tuning] adds trainable prompt embeddings to the input that is optimized by a prompt encoder to find a better prompt, eliminating the need to manually design prompts. The prompt tokens can be added anywhere in the input sequence, and P-Tuning also introduces anchor tokens for improving performance.
  • Prefix Tuning [LiL20] prefixes a series of task-specific vectors to the input sequence that can be learned while keeping the pretrained model frozen. The prefix parameters are inserted in all of the model layers.
  • IA3 [IA3] enhances efficiency by infusing learned vectors into transformer architectures, drastically cutting trainable parameters while preserving performance and minimizing inference latency.
  • AdaLoRA [Adalora] is a method for optimizing the number of trainable parameters to assign to weight matrices and layers, unlike LoRA, which distributes parameters evenly across all modules. More parameters are budgeted for important weight matrices and layers while less important ones receive fewer parameters.
Second, we extend the experiments exploring HydraLoRA on multiple datasets compared with more weighted average methods:
  • LoRA MoE [Zadouri2023] combines lightweight experts (LoRA) with MoE architecture for high efficiency, generalizing to new tasks without prior knowledge.
  • Lorahub [HuangLoraHub2023] employs black-box optimization to learn weights of 20 randomly selected LoRAs for new tasks, using weighted averaging without needing gradient calculations.
A detailed description of the baseline models can be found in Appendix A.
4.2 Overall Performance
Table 2: Comparative performance of different tuning schemes across multiple benchmarks on a single domain. 8-shot for GSM8K, zero-shot for others. # $\bar{B}$ refers to the average B matrix number of all datasets
Schemes MMLU Medical Law HumanEval GSM8K %Param # A # $\bar{B}$
P@1 P@10
LLaMA2-7B 38.88 35.98 33.51 13.10 20.34 10.38 - - -
@@ 4.2 Overall Performance
Prompt Tuning 39.91 37.59 35.02 13.66 21.55 13.18 0.001 - -
P-Tuning(256) 41.11 39.81 36.72 13.60 21.13 15.56 0.193 - -
(IA)3 40.45 37.12 35.25 13.54 23.17 13.98 0.009 - -
Prefix Tuning 41.78 40.28 36.54 13.23 22.56 16.89 0.077 - -
AdaLoRA(r=8) 44.32 42.83 39.36 14.81 23.78 19.51 0.093 1 1
LoRA(r=8) 43.22 41.59 37.85 15.67 22.95 18.24 0.062 1 1
@@
[1.5pt]
@@
We evaluated the Hydra structure LoRA — HydraLoRA that is characterized by a shared LoRA A matrix, while maintaining distinct B matrices that are trained separately. This configuration was meticulously compared with both the standard LoRA and the LoRA-Split approaches, emphasizing efficiency parameters. According to the results presented in Table 2, unlike LoRA-Split which straightforwardly adopts multiple vanilla LoRAs, HydraLoRA adopts an asymmetric LoRA structure that not only improves parameter efficiency by separating the uses of A matrix for commonalities and B matrices for diversities with a notably smaller adapter parameter set, but also employs a trainable router to improve the composition of multiple B matrices that outperforms the LoRA-Split approach. This finding is significant as it suggests that HydraLoRA not only enhances performance efficiency but also boosts overall system effectiveness. This may be driven by 1) different B modules capturing different features of the data-intrinsic knowledge, mitigating mutual interferences, and avoiding performance offsets. 2) Module A maintains the collaborative knowledge by taking the strengths of each and integrating them to improve the model performance.
While we hypothesize that the asymmetry is mainly rooted in the different initialization methods of A and B matrices, it is possible that this behavior varies on different model architectures and datasets. To the best of our ability, we extended the experiments exploring HydraLoRA on multiple datasets. LoRA MoE and their variants typically aim at tackling multi-tasks by employing multiple independent LoRAs. This makes them suitable for handling various domains. However, for a single dataset like ours, a “default” MoE method might not be optimal. HydraLoRA addresses this by constructing asymmetric structures and utilizing multiple B-modules to capture the specific nuances within the single dataset. The effectiveness of this approach is demonstrated by the experimental results in Table 3.
Table 4: Energy consumption and latency during fine-tuning with different LoRA approaches (fine-tuning LLaMA2-7B with GSM-8K).
Module Energy (KWh) Latency (h)
CPU GPU RAM
LoRA (r=8) 0.058 0.636 0.063 1.358
LoRA(r=16) 0.126 1.364 0.127 2.814
LoRA(r=32) 0.263 2.755 0.280 5.481
LoRA-Split(4×8) 0.272 2.908 0.291 5.633
HydraLoRA(r=8) 0.127 1.379 0.132 2.875
4.3 Energy and throughput analysis
We evaluate the system efficiency of HydraLoRA from two perspectives: training energy consumption and latency. The following experiments were executed on a GPU infrastructure consisting of 4 NVIDIA A40 GPUs and a CPU powered by an Intel(R) Xeon(R) Gold 6330 CPU clocked at 2.00GHz. Power consumption measurements were recorded using CodeCarbon [carbon]. Table 4 shows the results of various fine-tuning approaches for GSM-8K using the LLaMA2-7B model. we can see that HydraLoRA effectively speeds up the training process 1.96×1.96$\times$1.96 × and reduces 49.6% energy cost compared to LoRA (rank=32). While the energy consumption and latency of LoRA-Split exceeds the LoRA (rank=32). This is for the reason that HydraLoRA jointly considers inherent knowledge modularity and collaboration, which utilizes the “Hydra” structure with a shared A matrix and different B matrix. In this way, it only employs rank=16 training overhead but expands to a performance enhancement of more than rank=32. Overall, this experiment demonstrates the parameter effectiveness of HydraLoRA.
4.4 Ablation Study
To delve deeper into understanding the contributions of each component in HydraLoRA. we present the results of our ablation study in Table 5. The variant w/o MOE (essentially reverts to LoRA) excludes the MOE architecture. Similarly, the w/o gate variant employs uniform expert weights bypassing the gate function. The w/o hydra adopts multiple vanilla LoRAs in a straightforward way. Table 5 indicates that the full HydraLoRA model outperforms its variants, showing that both the MoE architecture and gate function significantly contribute to its effectiveness across various language understanding domains.
Table 5: Comparative performance of ablation study for HydraLoRA across multiple benchmarks.
Model MMLU Medical Law
w/o MoE 43.22 41.59 37.85
w/o gate 42.81 40.92 37.12
w/o hydra 46.94 45.28 41.35
HydraLoRA 47.22 45.71 42.18
Figure 5: The results of experiments for hyper-parameters K-value.
4.5 Hyper-parameter Analysis
As Figure 5 shown, we conduct a comprehensive and meticulous analysis by fine-tuning the Dolly-15K model on the LLaMA2-7B dataset and subsequently evaluating its performance on the MMLU benchmark to rigorously examine the impact of variations in the intrinsic component, symbolized by the variable $K$, on the model’s overall performance. Empirically we find that the K value is not a sensitive parameter for HydraLoRA, with a wide range of reasonable K values (e.g. 2 to 4) performing decently well in all settings in our experiments. Specifically, as Figure 5 (a) red line shown, the performance loss of K=3 vs. the optimal K=4 value is only 0.42%.
Therefore, based on this observation, we choose K-means because it is simple but effective, more sophisticated hyperparameter search approaches (e.g. parameter sweep or Bayesian optimization) will be unnecessarily costly. It’s noteworthy that HydraLoRA is adeptly designed to orchestrate its components in a way that it can automatically calibrate and navigate toward the optimal performance configuration across various parameters. This intelligent auto-tuning is achieved through the application of the k-means clustering algorithm. This strategic component orchestration not only enhances performance but also ensures a more efficient and effective utilization of resources, underpinning the model’s capability to adapt and perform efficiently in a dynamic computational environment.
5 Related work
Parameter-Efficient Fine-tuning
LLMs are becoming increasingly powerful, but fine-tuning them often requires significant computational resources. This has spurred research on parameter-efficient fine-tuning (PEFT) techniques that reduce memory and storage costs during adaptation. One prominent PEFT approach is adapters [HoulsbyGJMLGAG19, RebuffiBV17]. It introduces new, trainable dense layers within the existing model, keeping the original parameters frozen. This concept has proven successful across various domains [PfeifferKRCG21, Stickland019, Sung0B22, AutoPEFT]. Further improvements on adapter compactness involve constructing parameter matrices using Kronecker products of low-rank matrices [MahabadiHR21]. Another PEFT strategy directly manipulates activations with learned vectors. This can be achieved through concatenation [P-tuning, LiL20, LesterAC21], multiplication (IA3; [IA3]), or addition (BitFit; [ZakenGR22]). Prefix-tuning [LiL20] and prompt-tuning [LesterAC21] are noteworthy examples that fine-tune continuous prompts instead of designing discrete ones [DengWHWGSSXH22]. Interestingly, a study suggests that many PEFT methods can be viewed as a form of adapter, providing a unified perspective [HeZMBN22]. Beyond adding new parameters or altering the computational graph, researchers also explore sparse [GuoRK20, SungNR21] or low-rank updates (LoRA; [HuSWALWWC22]).
Multi-LoRA Architecture
LoRA has notably garnered increasing interest recently, becoming a standard approach for adapting LLMs such as LLaMA [llama, llama2] under limited computational resources. Recognizing its potential, researchers have delved deeper, exploring the benefits of employing multiple LoRAs. LoraHub [HuangLoraHub2023] takes this multi-LoRA approach by training several adapters and strategically picking combinations based on the domain during inference. Meanwhile, MultiLoRA [WangMultiLoRA2023] focuses on horizontal scaling, aiming to reduce LoRA’s parameter dependence. This involves splitting LoRA modules along the rank dimension and introducing learnable scaling factors for enhanced expressiveness. Addressing scaling challenges from a different angle, the mixture of LoRA concept is further proposed [Zadouri2023]. This mitigates resource consumption when scaling instruction-tuned LLMs. Recognizing the potential for conflict during instruction tuning, LoRAMoE [DouLoRAMoE2023] leverages the Mixture-of-Experts (MoE; [JacobsJNH91]) structure to safeguard the pre-trained LLM’s knowledge from excessive corruption by instruction data. Similarly, MOELoRA [LiuMOELoRA2023] incorporates a MoE framework into LLMs, thereby improving their multitasking capabilities in the medical domain. Shifting the focus to the system perspective, S-LoRA [ShengSLoRA2023] provides a framework for efficiently serving multiple LoRA adapters. Unlike previous methods that relied on choosing LoRA combinations based on their training domains, HydraLoRA breaks free from the dependence on domain knowledge during inference. Additionally, HydraLoRA’s asymmetric structure further enhances parameter efficiency compared to existing symmetric approaches.
6 Conclusion
In this work, we start by conducting exploratory experiments applying the LoRA technique to LLaMA2, aiming to adapt it to a new domain across various tasks. This study unveils the limitations of employing a single LoRA for the entire domain, highlighting the detrimental effects of domain interference. In response, we introduce a novel architecture HydraLoRA that features an asymmetric structure with a shared matrix for all samples and distinct matrices for each intrinsic component. This design improves domain adaptation by selectively focusing on distinct components, enhancing both fine-tuning and inference efficiency. Our research highlights the importance of balancing learning capabilities for diverse datasets against the need for a lean model, offering a viable pathway for improving LLMs with minimal parameter growth.
Appendix A Datasets and Baselines
A.1 Datasets
Single Domain
  1. 1.
  2. 2.
    Medical: we fine-tune with GenMedGPT and clinic-10k from ChatDoctor for medicine applications and evaluate medical tasks in MMLU including three related tasks: "clinical knowledge”, "professional medicine” and "college medicine”.
  3. 3.
    Law: we fine-tune with two legal instruction tuning datasets Lawyer-Instruct and US-Terms then evaluate with law tasks in MMLU including two related tasks: "professional law” and "international law”.
  4. 4.
  5. 5.
Multi-task Domain
We summarize the details of the used datasets as follows:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
A.2 Baselines
PEFT methods
  1. 1.
    Full Fine-tuning: is the default strategy for adaptation. During fine-tuning, the model is initialized with pretrained weights and biases, and all model parameters undergo gradient updates.
  2. 2.
    Prompt Tuning: adds task-specific prompts to the input, and these prompt parameters are updated independently of the pretrained model parameters which are frozen.
  3. 3.
    P-Tuning: adds trainable prompt embeddings to the input that is optimized by a prompt encoder to find a better prompt, eliminating the need to manually design prompts. The prompt tokens can be added anywhere in the input sequence, and P-Tuning also introduces anchor tokens for improving performance.
  4. 4.
    Prefix Tuning: prefixes a series of task-specific vectors to the input sequence that can be learned while keeping the pretrained model frozen. The prefix parameters are inserted in all of the model layers.
  5. 5.
  6. 6.
Multiple LoRA weighted average methods
  1. 1.
    $s_{i}=R(x)_{i}=\text{softmax}(W_{g}^{T}x)\quad(\text{Router})$ (6)
    $y=\sum_{i=1}^{n}s_{i}\cdot E_{i}(x)\quad(\text{MoE})$ (7)
  2. 2.
Appendix B LoRA Breakdown
Figure 6: Breakdown analysis of LoRA modules. Compare fine-tuned LoRA modules of GSM-8K [gsm8k] with its subsets using T-SNE. We employ the Independent and Identically Distributed (IID) segmentation scheme to divide GSM8K into three subsets and fine-tune them using different LoRAs. Consider LLaMA2-7b (random seed=42), which contains 32 decoder layers, corresponding to 32 adaptive modules. Each module consists of {0: q_proj_A, 1: q_proj_B, 2: v_proj_A, 3: v_proj_B} submodules. This makes a total of $32\times 4$ submodules. (a,b) left displays all submodules. (a,b) center shows all even submodules, i.e. the A-matrix. (a,b) right represents all odd submodules, i.e. the B-matrix. It can be seen that the differences in the fine-tuned LoRA modules for different tasks arise mainly from the B matrix.

### v2

@@ 3.2.2 Inference
During inference, HydraLoRA merges adapters by enabling routing computation based on the input. Specifically, since matrices B operate as linear functions, we initially compute a weighted average of the experts. Following this, we apply a PEFT transformation using the combined expertise. The HydraLoRA significantly enhances training efficiency through an extremely parameter-efficient MoE formulation. Additionally, the intrinsic structural modularity of HydraLoRA facilitates rapid recovery and merging of the trained parameters during inference, leading to substantial memory savings.
Table 2: Comparative performance of different tuning schemes across multiple benchmarks on a single domain. 8-shot for GSM8K, zero-shot for others. #$\bar{B}$ refers to the average $B$ matrix number.
HumanEval
Schemes MMLU Medical Law P@1 P@10 GSM8K %Param #A #$\bar{\textbf{B}}$
LLaMA2-7B [46] 38.88 35.98 33.51 13.10 20.34 10.38 - - -
@@
Prompt Tuning [23] 39.91 37.59 35.02 13.66 21.55 13.18 0.001 - -
P-Tuning(256) [28] 41.11 39.81 36.72 13.60 21.13 15.56 0.193 - -
Prefix Tuning [24] 41.78 40.28 36.54 13.23 22.56 16.89 0.077 - -
(IA)3 [26] 40.45 37.12 35.25 13.54 23.17 13.98 0.009 - -
AdaLoRA(r=8) [52] 44.32 42.83 39.36 14.81 23.78 19.51 0.093 1 1
LoRA(r=8) [17] 43.22 41.59 37.85 15.67 22.95 18.24 0.062 1 1
@@
4 Experiments
In this section, we detail the principal experiments. We begin with an overview of the experimental setup and implementation intricacies. Following this, we share our findings and offer a succinct interpretation.
4.1 Experiment Setting
Dataset and Benchmarks
To explore the properties and commonalities of the LoRA asymmetric structure, we conduct experiments on both single and multiple domains to evaluate the effectiveness of HydraLoRA for profiling intrinsic components. $\bullet$ Single domain. 1) General: we fine-tune with the general instruction tuning databricks-dolly-15k [8] for generic language capability and evaluate with MMLU [15]. 2) Medical: we fine-tune with GenMedGPT and clinic-10k from ChatDoctor [25] for medicine applications and evaluate medical tasks in MMLU. 3) Law: we fine-tune with two legal instruction tuning datasets Lawyer-Instruct [1] and US-Terms [4] then evaluate with law tasks in MMLU. 4) Math: we fine-tune with the training split of GSM8K [7] for mathematical reasoning and evaluate with test set of GSM8K. 5) Code: we fine-tune with CodeAlpaca [5] for code generation and evaluate with HumanEval [6]. $\bullet$ Multi-task domain. We select a portion of the Flanv2 [49] datasets covering Natural Language Understanding (NLU) and Natural Language Generation (NLG), which can be grouped into 10 distinct task clusters. Then we evaluate it with the Big-Bench Hard (BBH) [44] benchmark. A detailed description of the benchmarks can be found in Appendix A.1.
Baselines
$\bullet$ First, we compare HydraLoRA with different PEFT methods on single datasets: 1) Full fine-tuning; 2) Prompt Tuning [23]; 3) P-Tuning [28]; 4) Prefix Tuning [24]; 5) IA3 [26]; 6) AdaLoRA [52]. $\bullet$ Second, we extend the experiments exploring HydraLoRA on multiple datasets compared with more weighted average methods: 1) Lorahub [18] employs black-box optimization to learn weights of 20 randomly selected LoRAs for new tasks, using weighted averaging without needing gradient calculations. 2) LoRA MoE [50] combines lightweight experts (LoRA) with MoE architecture for high efficiency, generalizing to new tasks without prior knowledge. A detailed description of the baseline models can be found in Appendix A.2.
4.2 Overall Performance
Table 3: Comparative performance of different tuning schemes, including base model (Base), LoRA tuning (LoRA), LoraHub learning, multi-LoRA tuning with MoE inference (LoRA MoE) and our proposed HydraLoRA learning across mix-task domain on the BBH benchmark with LLaMA2-7B, LLaMA2-13B as the base LLM (3-shot). Refer to Appendix D for details.
Metrics Base [46] LoRA [17] Lorahub [18] LoRA MoE [50] HydraLoRA
Performance 7B 31.6 36.8 39.7 40.3 41.5
13B 38.4 40.1 41.9 43.7 44.1
# of A/B for training 0/0 1/1 48/48 48/48 1/10
# of A/B for inference 0/0 1/1 20/20 48/48 1/10
% Params - 0.062 1.240 2.976 0.341
We evaluated the Hydra structure LoRA — HydraLoRA that is characterized by a shared LoRA A matrix, while maintaining distinct B matrices that are trained separately. This configuration was meticulously compared with both the standard LoRA and the LoRA-Split approaches, emphasizing efficiency parameters.
According to the results presented in Table 2, unlike split which straightforwardly adopts multiple vanilla LoRAs, HydraLoRA adopts an asymmetric LoRA structure that not only improves parameter efficiency by separating the uses of A matrix for commonalities and B matrices for diversities with a notably smaller adapter parameter set, but also employs a trainable router to improve the composition of multiple B matrices that outperforms the LoRA-Split approach. This finding is significant as it suggests that HydraLoRA not only enhances performance efficiency but also boosts overall system effectiveness. This may be driven by 1) different B matrices capturing different features of the data-intrinsic knowledge, mitigating mutual interferences, and avoiding performance offsets. 2) Module A maintains the collaborative knowledge by taking the strengths of each and integrating them to improve the model performance.
While we hypothesize that the asymmetry is mainly rooted in the different initialization methods of A and B matrices, it is possible that this behavior varies on different model architectures and datasets. To the best of our ability, we extended the experiments exploring HydraLoRA on multiple datasets. LoRA MoE and their variants typically aim at tackling multi-tasks by employing multiple independent LoRAs. This makes them suitable for handling various domains. However, for a single dataset like ours, a “default” MoE method might not be optimal. HydraLoRA addresses this by constructing asymmetric structures and utilizing multiple B matrices to capture the specific nuances within the single dataset. The effectiveness of this approach is demonstrated by the experimental results in Table 3.
Refer to captionRefer to caption
Figure 5: Energy consumption and latency during fine-tuning with different LoRA approaches (fine-tuning LLaMA2-7B with GSM-8K).
4.3 Energy and Throughput Analysis
We evaluate the system efficiency of HydraLoRA from two perspectives: training energy consumption and latency. The following experiments were executed on a GPU infrastructure consisting of 4 NVIDIA A40 GPUs and a CPU powered by an Intel(R) Xeon(R) Gold 6330 CPU clocked at 2.00GHz. Power consumption measurements were recorded using CodeCarbon [33]. Figure 5 shows the results of various fine-tuning approaches for GSM-8K using the LLaMA2-7B model. we can see that HydraLoRA effectively speeds up the training process 1.96×1.96$\times$1.96 × and reduces 49.6% energy cost compared to LoRA (rank=32). While the energy consumption and latency of LoRA-Split exceeds the LoRA (rank=32). This is for the reason that HydraLoRA jointly considers inherent knowledge modularity and collaboration, which utilizes the “Hydra” structure with a shared A matrix and different B matrix. In this way, it only employs rank=16 training overhead but expands to a performance enhancement of more than rank=32. Overall, this experiment demonstrates the parameter effectiveness of HydraLoRA.
Figure 6: Comparative performance of ablation study for HydraLoRA across multiple benchmarks.
4.4 Ablation Study
To delve deeper into understanding the contributions of each component in HydraLoRA. we present the results of our ablation study in Figure 6. The variant w/o MoE (essentially reverts to LoRA) excludes the MoE architecture. Similarly, the w/o gate variant employs uniform expert weights bypassing the gate function. The w/o hydra adopts multiple vanilla LoRAs in a straightforward way. Figure 6 indicates that the full HydraLoRA model outperforms its variants, showing that both the MoE architecture and gate function significantly contribute to its effectiveness across various language understanding domains.
4.5 Hyper-parameter Analysis
As Figure 8 shown, we conduct a comprehensive and meticulous analysis by fine-tuning the Dolly-15K model on the LLaMA2-7B dataset and subsequently evaluating its performance on the MMLU benchmark to rigorously examine the impact of variations in the intrinsic component, symbolized by the variable $k$, on the model’s overall performance. Empirically we find that the number k of clusters is not a sensitive parameter for HydraLoRA, with a wide range of reasonable number k of clusters (e.g. 2 to 4) performing decently well in all settings in our experiments. Specifically, the performance loss of k=3 vs. the optimal k=4 is only 0.42%. Meanwhile, as illustrated in Figure 8, we employ three distinct methods to generate the number of corpus clusters 15-fold, and the results demonstrate that the k-means [29] yields comparable outcomes with DBSCAN [38]. Therefore, based on this observation, we choose k-means because it is simple but effective, more sophisticated hyperparameter search approaches (e.g. DBSCAN, parameter sweep and Bayesian optimization) will be unnecessarily costly. It’s noteworthy that HydraLoRA is adeptly designed to orchestrate its components in a way that it can automatically calibrate and navigate toward the optimal performance configuration across various parameters. This intelligent auto-tuning is achieved through the application of the k-means clustering algorithm. This strategic component orchestration not only enhances performance but also ensures a more efficient and effective utilization of resources, underpinning the model’s capability to adapt and perform efficiently in a dynamic computational environment.
Figure 7: Number of clusters generated by different approaches including developer-specific (static), k-means, and DBSCAN.
Figure 8: The results of experiments for hyper-parameters number of clusters.
5 Related work
Parameter-Efficient Fine-tuning
LLMs are becoming increasingly powerful, but fine-tuning them often requires significant computational resources. This has spurred research on parameter-efficient fine-tuning (PEFT) techniques that reduce memory and storage costs during adaptation. One prominent PEFT approach is adapters [16, 36]. It introduces new, trainable dense layers within the existing model, keeping the original parameters frozen. This concept has proven successful across various domains [34, 41, 42, 53]. Further improvements on adapter compactness involve constructing parameter matrices using Kronecker products of low-rank matrices [30]. Another PEFT strategy directly manipulates activations with learned vectors. This can be achieved through concatenation [28, 24, 23], multiplication (IA3; [26]), or addition (BitFit; [51]). Prefix-tuning [24] and prompt-tuning [23] are noteworthy examples that fine-tune continuous prompts instead of designing discrete ones [9]. Interestingly, a study suggests that many PEFT methods can be viewed as a form of adapter, providing a unified perspective [13]. Beyond adding new parameters or altering the computational graph, researchers also explore sparse [12, 43] or low-rank updates (LoRA; [17]).
Multi-LoRA Architecture
LoRA has notably garnered increasing interest recently, becoming a standard approach for adapting LLMs such as LLaMA [45, 46] under limited computational resources. Recognizing its potential, researchers have delved deeper, exploring the benefits of employing multiple LoRAs. LoraHub [18] takes this multi-LoRA approach by training several adapters and strategically picking combinations based on the domain during inference. Meanwhile, MultiLoRA [48] focuses on horizontal scaling, aiming to reduce LoRA’s parameter dependence. This involves splitting LoRA modules along the rank dimension and introducing learnable scaling factors for enhanced expressiveness. Addressing scaling challenges from a different angle, the mixture of LoRA concept is further proposed [50]. This mitigates resource consumption when scaling instruction-tuned LLMs. Recognizing the potential for conflict during instruction tuning, LoRAMoE [11] leverages the Mixture-of-Experts (MoE; [19]) structure to safeguard the pre-trained LLM’s knowledge from excessive corruption by instruction data. Similarly, MOELoRA [27] incorporates a MoE framework into LLMs, thereby improving their multitasking capabilities in the medical domain. Shifting the focus to the system perspective, S-LoRA [40] provides a framework for efficiently serving multiple LoRA adapters. Unlike previous methods that relied on choosing LoRA combinations based on their training domains, HydraLoRA breaks free from the dependence on domain knowledge during inference. Additionally, HydraLoRA’s asymmetric structure further enhances parameter efficiency compared to existing symmetric approaches.
6 Conclusion
In this work, we start by conducting exploratory experiments applying the LoRA technique to LLaMA2, aiming to adapt it to a new domain across various tasks. This study unveils the limitations of employing a single LoRA for the entire domain, highlighting the detrimental effects of domain interference. In response, we introduce a novel architecture HydraLoRA that features an asymmetric structure with a shared matrix for all samples and distinct matrices for each intrinsic component. This design improves domain adaptation by selectively focusing on distinct components, enhancing both fine-tuning and inference efficiency. Our research highlights the importance of balancing learning capabilities for diverse datasets against the need for a lean model, offering a viable pathway for improving LLMs with minimal parameter growth. More discussion about limitation and broader impacts are available in Appendix E and F.
Appendix A Datasets and Baselines
A.1 Datasets
Single Domain
  1. 1.
  2. 2.
    Medical: we fine-tune with GenMedGPT and clinic-10k from ChatDoctor for medicine applications and evaluate medical tasks in MMLU including three related tasks: “clinical knowledge”, “professional medicine” and “college medicine”.
  3. 3.
    Law: we fine-tune with two legal instruction tuning datasets Lawyer-Instruct and US-Terms then evaluate with law tasks in MMLU including two related tasks: “professional law” and “international law”.
  4. 4.
  5. 5.
Multi-task Domain
We summarize the details of the used datasets as follows:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
A.2 Baselines
PEFT methods
  1. 1.
  2. 2.
    Prompt Tuning adds task-specific prompts to the input, and these prompt parameters are updated independently of the pretrained model parameters which are frozen.
  3. 3.
    P-Tuning adds trainable prompt embeddings to the input that is optimized by a prompt encoder to find a better prompt, eliminating the need to manually design prompts. The prompt tokens can be added anywhere in the input sequence, and P-Tuning also introduces anchor tokens for improving performance.
  4. 4.
    Prefix Tuning prefixes a series of task-specific vectors to the input sequence that can be learned while keeping the pretrained model frozen. The prefix parameters are inserted in all of the model layers.
  5. 5.
  6. 6.
Multiple LoRA weighted average methods
  1. 1.
    $s_{i}=R(x)_{i}=\text{softmax}(W_{g}^{T}x)\quad(\text{Router})$ (5)
    $y=\sum_{i=1}^{n}s_{i}\cdot E_{i}(x)\quad(\text{MoE})$ (6)
  2. 2.
Appendix B Initialization via k-means
In the case of considering heterogeneous corpora, it is crucial to select the appropriate number $N$ of matrix B, to ensure consistent performance and minimize unnecessary computational overhead. This choice is usually closely related to the training corpus. In this work, we propose initializing HydraLoRA modules via $k$-means [29] algorithm for adaptive initialization. Specifically, $k$-means is utilized to process the heterogeneous corpus to identify the best-fit taxonomy of the corpus, i.e., the optimal $k$. First, we extract key features from the corpus by applying the Term Frequency-Inverse Document Frequency (TF-IDF; [37]) algorithm and transform the textual information into numerical feature vectors. We integrate the elbow method [22] to determine the optimal value of $K$. Initially, $K$ cluster centers are randomly selected for preliminary clustering as Eq. 7, followed by updating the cluster centers to accurately reflect the data within each cluster as Eq. 8. where $C_{j}$ is the cluster center to which data point $X_{i}$ is assigned and $d(\cdot,\cdot)$ is the Euclidean distance function. $S_{j}$ is the set of data points in the $j$-th cluster.
$C_{j}$ $=\underset{C_{j}}{\mathrm{argmin}}\ d(X_{i},C_{j})$ (7)
$C_{j}$ $=\frac{1}{|S_{j}|}\sum_{X_{i}\in S_{j}}X_{i}$ (8)
Appendix C LoRA Breakdown
Refer to captionRefer to captionRefer to caption
(a) Compare fine-tuned LoRA modules of GSM8K [7] with its subsets using T-SNE. We employ the Independent and Identically Distributed (IID) segmentation scheme to divide GSM8K into three subsets and fine-tune them using different LoRAs.
Refer to captionRefer to captionRefer to caption
(b) Specific tasks.
Figure 9: Breakdown analysis of LoRA modules. Compare fine-tuned LoRA modules of GSM-8K [7] with its subsets using T-SNE. We employ the Independent and Identically Distributed (IID) segmentation scheme to divide GSM8K into three subsets and fine-tune them using different LoRAs. Consider LLaMA2-7B (random seed=42), which contains 32 decoder layers, corresponding to 32 adaptive modules. Each module consists of {0: q_proj_A, 1: q_proj_B, 2: v_proj_A, 3: v_proj_B} submodules. This makes a total of $32\times 4$ submodules. (a,b) left displays all submodules. (a,b) center shows all even submodules, i.e. the A-matrix. (a,b) right represents all odd submodules, i.e. the B-matrix. It can be seen that the differences in the fine-tuned LoRA modules for different tasks arise mainly from the B matrix.
Appendix D More Results
Tabel  4 shows Comparative performance of different tuning schemes, including base model (Base), LoRA tuning (LoRA), LoraHub learning, multi-LoRA tuning with MoE inference (LoRA MoE) and our proposed HydraLoRA learning across mix-task domain on the BBH benchmark with LLaMA2-7B as the base LLM (3-shot).
@@ Appendix D More Results
Appendix E Limitation
HydraLoRA is computationally demanding, primarily due to the necessity of fine-tuning large-scale language models. It incurs a higher training expenditure than conventional PEFT methods, attributed to the employment of multiple adapter copies. Empirical data suggest that HydraLoRA requires 1 to 2 times more training iterations compared to typical PEFT methods, which adversely affects the environmental footprint of model training. The HydraLoRA framework, distinct from prevailing PEFT approaches, holds promise for enhancing the efficacy of existing PEFT strategies. In our current study, we examine established PEFT techniques — LoRA. However, we have not tested additional configurations such as prompt-tuning and adapter, deferring these explorations to subsequent research. Additionally, our assessment is exclusively within the context of fine-tuning. Exploration of its efficacy during the pre-training phase remains an avenue for future research.
Appendix F Broader Impacts
Positive Societal Impacts
The proposed HydraLoRA framework, with its asymmetric structure and parameter-efficient fine-tuning approach, has the potential to make LLMs more accessible and efficient. This could democratize AI, enabling more researchers, developers, and organizations to leverage the power of LLMs for various applications, ultimately driving innovation and progress. Moreover, by effectively addressing the challenge of domain or task interference, HydraLoRA could significantly enhance the performance of LLMs in complex, multi-task domains. This could lead to more accurate and reliable AI-powered tools and services in areas like healthcare, education, and finance, ultimately improving the quality of life for many people. Lastly, the parameter-efficient approach of HydraLoRA could help reduce the computational resources required for training and fine-tuning LLMs, thereby lessening the environmental impact of AI.
Negative Societal Impacts
As with any AI technology, there are potential negative societal impacts to consider. The risk of misuse is a significant concern, as HydraLoRA could be used for malicious purposes, such as creating more sophisticated and convincing deepfakes or spreading misinformation and propaganda. Additionally, the increased efficiency and accessibility of AI brought about by HydraLoRA could lead to job displacement in certain sectors, as AI-powered tools and services become capable of performing tasks traditionally done by humans. Lastly, the use of LLMs in various applications could potentially lead to privacy and security issues, especially if these models are used to process or generate sensitive information. The proposed HydraLoRA framework, while not directly related to these issues, could inadvertently contribute to them by making it easier to deploy LLMs in various applications.